OpenStudy (anonymous):

Please help, not quite getting it. 2cosα=√3 I know that you divide both sides by two and then you have cosα=√3/2 which is a ratio for a 30 60 90 but i dont know where to go from there.

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

is the answer the angle they share? the 30?

OpenStudy (anonymous):

yes, it is the 30 degree angle

OpenStudy (anonymous):

you've been like my best friend tonight, i feel like i should send you a gift basket...

OpenStudy (anonymous):

so it would be both the -30 and the +30...

OpenStudy (anonymous):



OpenStudy (anonymous):

if you are using the unit circle, take a look at the last page of the attached cheat sheet

OpenStudy (anonymous):

find the point on the unit circle where the first coordinate is \(\frac{\sqrt{3}}{2}\) and look at the angle it will be 30 degrees or if you are using radians it will be \(\frac{\pi}{6}\)

OpenStudy (anonymous):

okay, my answer though includes the -30

OpenStudy (anonymous):

yes that is also true there are an infinite number of answers

OpenStudy (anonymous):

you don't see -30 on the cheat sheet, but you can visualize going 30 degrees clockwise and you get the same first coordinate

OpenStudy (anonymous):

in math we say "cosine is even" meaning \(\cos(x)=\cos(-x)\) so for example \[\cos(30)=\cos(-30)\]

OpenStudy (anonymous):

okay that makes sense! Thank you so much for your help you have no idea how much of a blessing your explanations have been!

OpenStudy (anonymous):

please put a bottle of whiskey in the basket. thanks

OpenStudy (anonymous):

(yw)

OpenStudy (anonymous):

btw if you haven't seen it before, hold on to that cheat sheet, or if you lose it remember "paul's notes" you can google and find it there other good cheat sheets too

OpenStudy (anonymous):

:) No problem, it'll be in there! sweet! thanks :) I think this is the last you'll have to deal with me tonight though, I'm off to sleep. Thanks again, you're awesome.

OpenStudy (anonymous):

(blush) enough math for me too time for tv

OpenStudy (anonymous):

that sounds rather fantastic. Maybe I'll watch an episode of Suits before I check out...